A driver leaves on a road trip at 8:00am. The function v(t) is the speed he is traveling t hours after the start of the trip. What value of t should be used so that v(t) will represent the driver's speed at 9:40 am? (three decimal places)
Found 2 solutions by nerdybill, richard1234:
Answer by nerdybill(7384)

You forgot to include the expression for v(t)

Answer by richard1234(7193)

At 9:40, t = 1 hour 40 minutes = 5/3 hours, so use t = 5/3 or 1.667.